# Linying Lv

### v1

@@ 2.2 Information Content
I use out-of-sample $R^{2}$ as a measure for the information content of analyst reports. To project contemporaneous stock return on analyst information, I train Ridge regression models with the following equations each year with an expanding training window:
@@
$R_{\mathrm{oos}}^{2}=1-\frac{\sum_{(i,t)\in\mathcal{T}}\left(CAR_{[-1,+1],it}-\widehat{CAR}_{[-1,+1],it}\right)^{2}}{\sum CAR_{[-1,+1],it}^{2}},$ (3)
where $\mathcal{T}$ is the test set of $(i,t)$ which has not been used in training and validation of the model. $R_{\mathrm{oos}}^{2}$ aggregates estimation across analyst reports and captures the portion of return variance that can be explained by analysts’ information.
@@ 2.3 Information Value
I measure the value of information using equation (4) following Kadan and Manela (forthcoming) who show that it approximates the value of information in a strategic setting with multiple informed investors, each possessing an imperfect signal based on the theoretical framework of Back et al. (2000):
@@
where $\widehat{\Omega}_{it}$ is the ratio of explainable realized return variance estimated using Ridge regression to price impact estimated from 1-minute log returns and order flow, multiplied by the closing price of the day prior to the [t-1, t+1] window. [^3] Due to risk neutrality, the equilibrium of the continuous-time Kyle model holds if $\tilde{v}$ is an unbiased signal of the asset price. The intuition is illustrated with an extended single-auction Kyle model incorporating incomplete information in the Online Appendix D.
Specifically, $\widehat{r}_{ijt}$ is the estimated cumulative abnormal return $\widehat{CAR}_{[-1,+1]}$ of analyst report $j$ about stock $i$ issued on day $t$, and $r_{it}$ is the realized $CAR_{[-1,+1]}$, with abnormal returns calculated as the raw return minus the buy-and-hold return on the NYSE/AMEX/NASDAQ value-weighted market return. $N$ denotes the number of analyst reports about stock $i$ on day $t$. To estimate the price impact $\lambda_{it}$, I perform a regression of 1-minute log returns ($r_{itk}$) on contemporaneous share order flow ($y_{itk}$). Here, $y_{itk}$ is calculated as the change in cumulative signed order flows $Y_{it\tau_{k}}$ over intraday intervals. The signed order flow is classified as buys (+1) or sells (-1) based on algorithms from Lee and Ready (1991). [^4] The algorithms from Ellis et al. (2000) and Chakrabarty et al. (2007) are also used in robustness tests. The regression model used is
@@
where $\widehat{\lambda}_{it}$ represents the estimated price impact (or Kyle’s lambda). In this context, $r_{itk}=p_{it\tau_{k}}-p_{it\tau_{k-1}}$ and $y_{itk}=Y_{it\tau_{k}}-Y_{it\tau_{k-1}}$, with $p_{it\tau_{k}}$ denoting the log price of stock $i$ within the window [t-1, t+1] at time $\tau\in[0,T]$. The interval $T$ is divided into one-minute sections $\tau_{0},\tau_{1},\ldots,\tau_{K}$. Thus, each regression has $1170(=390*3)$ observations.
I employ the measure for two key reasons. First, the framework is rooted in a general setting of imperfect competition among informed traders, which closely aligns with the reality of financial markets where multiple investors have access to analyst information. This setting is particularly apt for analyzing the value of analyst reports, as it accounts for the correlated nature of information and the strategic interactions among market participants. Second, this measure provides a conservative lower bound for the value estimation, allowing quantification of the minimum potential value of analyst information without overstating its importance.
Given that the measure is a ratio of two random variables measured with error, interpretation for information value estimation is inherently nuanced. To tackle this issue, I apply the delta method in estimating the mean and variance of information value over sample $s$. The delta method used to derive the first-order estimator is discussed in the Online Appendix E.
@@ 2.4 Shapley Value Decomposation
To assess the impact of topics within analyst reports, I employ the Shapley Additive exPlanations (SHAP) methodology, as detailed by Lundberg and Lee (2017) and applied to financial contexts by Chen et al. (2022). SHAP is a solution concept proposed by Shapley (1953) in cooperative game theory that distributes the total payoff generated by a coalition of players among the players themselves. In my context, this allows me to attribute the overall explanatory power of an analyst report to its constituent topics. Leveraging the additive nature of text embeddings, I design a novel decomposition method based on SHAP values. This approach enables me to systematically assess the relative importance of various topics in analyst reports.
First, LLMs pool report-level embeddings by averaging embeddings across all tokens and all layers. The textual embedding can be represented as:
@@
One concern in the architecture of a transformer model is that each token in the input sequence can pay attention to other words through a mechanism called “self-attention” or “scaled dot-product attention.” This mechanism allows the model to weigh the importance of each word relative to every other word in the sequence, thereby capturing contextual information. Consequently, the embedding of a word contains information from other words, which may pertain to different topics. The embedding of the word might predominantly reflect the context information from other topics. To isolate the self-attention across topics, I process the embedding for each sentence in analyst reports and take the token-weighted mean as the report embedding.
@@
The loss in $R_{\mathrm{oos}}^{2}$ from full-context embedding to sentence-segmented embedding reflects the information content of context in analyst reports.
Second, I calculate the Shapley value of the sentence-segmented embeddings distributed to each topic. Empirically, the Shapley value can be written as:
@@
where the sum extends over all subsets $S$ of $P$ not containing topic $p$, including the empty set. Note that $\binom{n}{k}$ is the binomial coefficient. $y_{p}^{\text{emb}}$ is the token-weighted sentence embedding of topic $p$, and $y_{s}^{\text{emb}}$ is the token-weighted sentence embedding of all topics in subset $S$. For each combination of topic embeddings, I calculate the $R_{\text{oos}}^{2}$ accordingly. [^5] Shapley value requires an input for $X$ when it is missing. I assign a value of zero, treating the embedding as having all dimensions set to zero for model estimation. This approach fairly distributes the total $R^{2}_{\mathrm{OOS}}$ among topics, revealing their relative importance in explaining market reactions to analyst reports.
@@ 3 Data and Empirical Results
This section presents the empirical analysis of analyst report information. I begin by examining the information content of these reports and decomposing it across different topics. I then quantify the strategic dollar value of analyst information. Cross-sectional analyses and robustness checks are conducted to validate the main findings.
@@ 3.1 Data
Sell-side analyst reports are obtained from the Mergent Investext Database. I download 223,091 analyst reports of S&P 100 firms from 2000 to 2023. The initial reports are in PDF format, and I use Adobe Acrobat to extract structured text from each report. The segmented text is chunked into sentences. To get clean content of analyst discussions, I train a BERT model to remove the boilerplate segments from each report. [^6] Boilerplate segments disclose information about the brokerage firm and analyst research team and do not reflect analyst opinions. Following Li et al. (2023), I randomly sampled one report from each of the top 20 brokerage firms producing the largest number of analyst reports and manually labeled each sentence in those reports. The BERT model trained for boilerplate classification reached an accuracy score of 97.31% on a testing sample.
@@
The Mergent Investext Database provides information on analyst names and the release date for each report. To connect Investext with the I/B/E/S database, I first match the lead analyst name in reports to Analyst ID (AMASKCD) following the procedure of Li et al. (2023). I then match specific reports to the corresponding EPS forecasts, price targets, and recommendation files in the I/B/E/S database. Following Huang et al. (2014), I define the matching window for each report as 2 days before the announcement date to 2 days after the revision date, accounting for multiple reports released between these dates. [^7] In I/B/E/S, each analyst forecast for a specific firm has two time stamps. The announcement date (ANNDATS) is the date the analyst announces the new forecast and revises the old one. The revision date (REVDATS) is the date the forecast becomes invalid or is revised. Reports are linked to I/B/E/S forecasts by matching analyst names, firm tickers, and this time window. The numerical measures include earnings forecasts, stock recommendations, and price target information from analyst reports. I also compile firm characteristics from the CRSP and Compustat databases. [^8] See Table A2 and Table A3 for detailed definitions and summary statistics. After excluding analyst reports released on earnings announcement days, my final sample comprises 122,252 reports of S&P 100 firms from 1,305 analysts across 140 unique brokerage firms.
@@
Table 1 presents the summary statistics of the analyst report data, including the number of reports, the number of unique brokerage firms, and the number of unique analysts for each year, or across the Fama-French 12 industries. It also reports the average number of pages and tokens per report, which are important indicators of report lengths. Overall, the time trend in the number of analyst reports aligns with that reported in Bonini et al. (2023), and the average number of pages per report is consistent with that in Huang et al. (2014). [^9] The data reveals a peak in the number of analyst reports in 2013, followed by a subtle decline. This trend can be attributed to the regulatory impacts of the Dodd-Frank Wall Street Reform in the United States and the introduction of MiFID II in Europe (Bradshaw et al., 2017).
@@
The daily stock-level price impact is estimated using intraday data from the NYSE TAQ database, covering an out-of-sample period from January 2, 2015, to December 28, 2023. To ensure comparability over time, all prices are adjusted for inflation using the Consumer Price Index (CPI), with December 2020 as the base period.
@@
In Section 3.2.2, I explore textual information from earnings conference calls and subsequent analyst reports. The earnings call transcripts are sourced from the Seeking Alpha website, which has been collecting the transcripts since 2005101010Available at https://seekingalpha.com/earnings/earnings-call-transcripts. These transcripts provide comprehensive accounts of the calls, including executive presentations and Q&A sessions with analysts, offering valuable insights into manager-analyst interactions. I sampled earnings call transcripts of S&P 100 constituent firms from 2005 to 2023, merging them with ticker and Report Date for Quarterly Earnings (RDQ) data from Compustat. The matching window between Seeking Alpha’s publishing date and RDQ is set to two weeks. This process yielded a final sample of 7,909 firm-quarter-level conference calls.
Given my goal to estimate the information value of analyst outputs, I need to define the information set available to investors. The numerical information and variables used as inputs to machine learning models are as follows:
@@
Firm-Level Measures: $Size$ is measured by market equity. The Book-to-Market Ratio ($BtoM$) compares the firm’s book value to its market value. Following Huang et al. (2018), the Earnings Suprise ($SUE$) is calculated as the actual earnings per share (EPS) minus the last consensus EPS forecast prior to the earnings announcement date (EAD), with $AbsSUE$ representing its absolute value. The $Miss$ variable is a binary indicator set to one if $SUE$ is positive. $TradingVolume$ is the stock’s trading volume on the latest earnings announcement days standardized by the number of shares outstanding. The Distance to Default ($DD$) variable is sourced from the NUS Credit Research Initiative (CRI). $Fluidity$ measures the firm’s product market fluidity.
@@ 3.2 Information Content of Analyst Reports
In this section, I provide yearly $R^{2}_{\mathrm{OOS}}$ of analyst reports for explaining contemporaneous cumulative abnormal returns to examine the information content of analysts’ quantitative and qualitative outputs. Furthermore, I attribute the information content of analyst report text across topics using Shapley value decomposition.
@@ 3.2.1 Information Content of Quantitative and Qualitative Information
Table 2 compares the information content embedded in numerical and textual measures, quantified through their $R^{2}_{\mathrm{OOS}}$. The estimation model used in this table is the Ridge regression, where the objective variable is $CAR_{[-1,+1]}$ centered at the analyst report release date. For reports released on non-trading days, I adjust t=0 to the next trading day as recorded in the CRSP database. The estimation methodology employs an expanding window approach, constructing training sets from 2000 to the year preceding each target year. I estimate separate Ridge regression models for each input configuration, conducting yearly out-of-sample estimations of $\widehat{CAR}_{[-1,+1]}$, allowing for dynamic assessment of model performance over time.
@@
Panel A of the table delineates the performance of four distinct input configurations: (1) forecast revisions, encompassing analyst revisions on recommendations, earnings forecasts, and target prices; (2) comprehensive numerical measures, incorporating report-level, firm-level, industry-level, and macroeconomic indicators as detailed in Section 3.1; (3) textual embeddings, the 5,120-dimensional full-context representations of each report generated by LLaMA-2-13B model; and (4) a hybrid input combining revision measures with textual embeddings.
@@
Column (1) reveals an $R^{2}_{\mathrm{OOS}}$ of 9.01% for the revision input, corroborating extant literature (Brav and Lehavy, 2003; Asquith et al., 2005; Bradshaw et al., 2013) on the price impact of analyst forecast revisions. The inclusion of additional numerical variables in column (2) yields only a marginal improvement, with the $R^{2}_{\mathrm{OOS}}$ reaching 9.08% using 17 numerical features. I compare the estimation with a zero benchmark and report the Diebold-Mariano test t-statistics in columns (2) and (4). The Diebold-Mariano test follows a standard normal distribution under the null hypothesis of equal predictive accuracy between two competing forecasting models, allowing similar interpretations of t-statistics as in regression analysis (Gu et al., 2020). The models incorporating revision and numerical inputs significantly outperform a zero-estimation benchmark.
@@
Columns (5) and (6) elucidate the performance of text embedding input, showcasing a substantial $R^{2}_{\mathrm{OOS}}$ of 10.19% and a t-statistic of 8.20. This underscores the significant information content of the report text in explaining contemporaneous stock returns. The $R^{2}_{\mathrm{OOS}}$ exhibits relative stability across 2015-2023, with a notable exception during the 2020 pandemic recession. Sarkar and Vafa (2024) raise the look-ahead bias concern associated with pre-trained LLMs. Language model outputs may incorporate future information that should not be available. Since the LLMs I applied were trained on language data ending before 2023, the $R^{2}_{\mathrm{OOS}}$ in 2023 reflects true out-of-sample performance, confirming the robustness of the main findings during this period. The Diebold-Mariano test comparing revision-only and text-only inputs yields a t-statistic of 1.66, highlighting the substantial information content in the report text.
@@
In columns (7) and (8), I further investigate the performance of combining both revision and text input in the Ridge regression. The combination achieves an $R^{2}_{\mathrm{OOS}}$ of 12.28%, which represents a 3.27% improvement over the $R^{2}_{\mathrm{OOS}}$ of the revision-only model and a 2.09% improvement over the $R^{2}_{\mathrm{OOS}}$ of the text-only model. I report the DM t-statistic of the comparison between combined input with revision input and text input in columns (7)-(1) and (7)-(5). The overall t-statistics of 3.95 and 3.77 indicate a significant superiority in the combination of text data and revision measures in explaining contemporaneous $CAR$. The DM tests confirm that report text and forecast revision contain distinct information sets that are valued by market investors.
@@
To address the concern that the explanatory power might be derived from quantitative information in analyst reports, I conduct robustness checks using content with numbers removed. In Table A4, I exclude numerical data from the written reports before generating text embeddings. The $R^{2}_{\mathrm{OOS}}$ of the number-free text embedding input model is 10.95%, still significantly higher than that of quantitative revision measures (t-statistic = 2.40). This suggests that the superior performance of the text-based model is not driven by numerical information within the reports.
@@
LLMs are known to exhibit improved language skills as the size of their parameters increases. I provide a quantitative comparison of LLMs with varying numbers of parameters, including the BERT model, OpenAI GPT-3 embedding model, and LLaMA-3 model. The LLaMA models have a substantially larger number of parameters compared to the BERT and OpenAI models. Panel B reports the $R^{2}_{\mathrm{OOS}}$ and DM t-stats of pairwise comparisons between each model and the LLaMA-2 model. The input in Panel B is report text, comparable to column (3) in Panel A. A positive t-statistic indicates that the model outperforms the LLaMA-2 benchmark. Overall, the BERT and OpenAI models underperform the LLaMA-2 model at the 1% level for each year. The LLaMA-3 model gives an $R^{2}_{\mathrm{OOS}}$ of 9.66%, which is slightly better but statistically insignificant compared to the LLaMA-2 model. The results in Table 2 suggest that the out-of-sample information content of analyst report text is fairly robust across different LLMs. The pairwise comparisons between LLMs also shed light on how the scale of the model influences its performance capabilities.
@@
To test the robustness of the information content in these contemporaneous measures, I also use $CAR$ measures in 2-day and 5-day windows, as applied in Huang et al. (2014) and Asquith et al. (2005), respectively. Additionally, I examine the abnormal stock return on the days that reports are released. In Table A5, I find a similar magnitude in $R^{2}_{\mathrm{OOS}}$ for explaining $AR_{[0]}$, $CAR_{[0,+1]}$, and $CAR_{[-2,+2]}$, ranging from 6.92% to 7.57% for the out-of-sample period from 2015 to 2023. In Figure A3, I plot the information value of analysts’ information by industry, revealing a pattern consistent with the information content results.
@@
In Table A6, I present the information content of analyst report text across the Fama-French 12 industries. The $R^{2}_{\mathrm{OOS}}$ exceeds 10% in five industries: Shops, Other, Manufacturing, Chemicals, and Durables. Notably, the $R^{2}_{\mathrm{OOS}}$ and DM t-statistics are significantly lower in the Non-Durables, Telecommunications, Energy, and Utilities sectors. The industries with higher $R^{2}_{\mathrm{OOS}}$ typically involve complex products and services that benefit from detailed analyst insights, whereas those with lower $R^{2}_{\mathrm{OOS}}$ may operate in more stable or regulated environments where analyst reports add less incremental value.
To assess economic magnitude, I perform OLS regressions as specified in Equation (10):

### v2

@@ 2.2 Information Content
Following Asquith et al. (2005), I use $R^{2}$ as a measure for information content. To project contemporaneous stock return on analyst information, I train Ridge regression models with the following equations each year with an expanding training window:
@@
$R_{\mathrm{oos}}^{2}=1-\frac{\sum_{(i,t)\in\mathcal{T}}\left(CAR_{[-1,+1],it}-\widehat{CAR}_{[-1,+1],it}\right)^{2}}{\sum_{(i,t)\in\mathcal{T}}CAR_{[-1,+1],it}^{2}},$ (3)
where $\mathcal{T}$ is the test set of $(i,t)$ which has not been used in training and validation. $R_{\mathrm{oos}}^{2}$ aggregates estimation across analyst reports and captures the portion of return variance that can be explained by analysts’ information.
@@ 2.3 Information Value
I measure the value of information using equation (4) following Kadan and Manela (forthcoming) who show that it approximates the value of information in a strategic setting with multiple informed investors, each possessing an imperfect signal:
@@
where $\widehat{\Omega}_{it}$ is the ratio of explainable realized return variance estimated using Ridge regression to price impact estimated from 1-minute log returns and order flow, and $p_{it_{-}}$ is the closing price of day t-2. $\widehat{r}_{ijt}$ is the estimated cumulative abnormal return $\widehat{CAR}_{[-1,+1]}$ of analyst report $j$ about stock $i$ issued on day $t$, and $r_{it}$ is the realized $CAR_{[-1,+1]}$, with abnormal returns calculated as the raw return minus the buy-and-hold return on the NYSE/AMEX/NASDAQ value-weighted market return. $N$ denotes the number of analyst reports about stock $i$ on day $t$.
To estimate the price impact $\widehat{\lambda}_{it}$, I perform a regression of 1-minute log returns ($r_{itk}$) on contemporaneous share order flow ($y_{itk}$). Here, $y_{itk}$ is calculated as the change in cumulative signed order flows $Y_{it\tau_{k}}$ over intraday intervals. The signed order flow is classified as buys (+1) or sells (-1) based on algorithms from Lee and Ready (1991). [^2] The algorithms from Ellis et al. (2000) and Chakrabarty et al. (2007) are also used in robustness tests. The regression model is
@@
where $\widehat{\lambda}_{it}$ represents how sensitive price is to trading volume (or Kyle’s lambda). In this regression, $r_{itk}=p_{it\tau_{k}}-p_{it\tau_{k-1}}$ and $y_{itk}=Y_{it\tau_{k}}-Y_{it\tau_{k-1}}$, with $p_{it\tau_{k}}$ denoting the log price of stock $i$ within the window [t-1, t+1] at time $\tau\in[0,T]$. The 3-day interval $T$ is divided into one-minute sections $\tau_{0},\tau_{1},\ldots,\tau_{K}$. Thus, each regression has $1170(=390*3)$ observations.
Essentially, the measure reflects the ratio of return volatility to price impact. It is derived based on the theoretical framework of Back et al. (2000), which is a continuous-time version of Kyle (1985) model. Due to risk neutrality, the equilibrium holds if $\tilde{v}$ is an unbiased signal of the asset price. The intuition is illustrated with an extended single-auction Kyle model incorporating incomplete information in the Online Appendix D.
I employ the measure for two key reasons. First, the framework is rooted in a general setting of imperfect competition among informed traders, which closely aligns with the reality of financial markets where multiple investors have access to analyst information. Second, this measure provides a conservative lower bound for the value estimation, allowing quantification of the minimum potential value of analyst information without overstating its importance. [^3] Kadan and Manela (forthcoming) argue the measure also serves as an approximate upper bound when informed traders are uncorrelated, which does not apply to the analyst case.
Given that the measure is a ratio of two random variables measured with error, interpretation for information value estimation is inherently nuanced. To tackle this issue, I apply the delta method to estimate the mean and variance of information value over sample $s$. The delta method used to derive the first-order estimator is discussed in the Online Appendix E.
@@ 2.4 Shapley Value Decomposation
To assess the impact of topics within analyst reports, I employ the Shapley Additive exPlanations (SHAP) methodology. SHAP is a solution concept proposed by Shapley (1953) in cooperative game theory that distributes the total payoff generated by a coalition of players among the players themselves. It was later applied for machine learning model interpretation by Lundberg and Lee (2017) and Chen et al. (2022). In my context, the approach allows me to attribute the overall explanatory power of an analyst report to its constituent topics. Leveraging the additive nature of text embeddings, I design a novel decomposition method to systematically assess the relative importance of various topics in analyst reports.
LLMs pool report-level embeddings by averaging embeddings across all tokens and all layers. The textual embedding can be represented as:
@@
One concern raised by the nature of transformer models is that each token in the input sequence can pay attention to other tokens through a mechanism called “self-attention” or “scaled dot-product attention.” This architecture allows the model to weigh the importance of each token relative to every other token and capture contextual information. For example, the embedding of token $i$ contains information from token $j$, which may pertain to a different topic. To isolate the self-attention across topics, I process the embeddings sentence by sentence to ensure the token embeddings only reflect the context information within the sentence. I then take the token-weighted mean of sentence embeddings as the report representation, which is referred to as sentence-segment embeddings.
@@
The loss in $R_{\mathrm{oos}}^{2}$ from full-context embedding to sentence-segmented embedding reflects the information content of context across sentences.
Second, I calculate the Shapley value of the sentence-segmented embeddings distributed to each topic. Specifically, the Shapley value can be written as:
@@
where the sum extends over all subsets $S$ of $P$ not containing topic $p$, including the empty set. Note that $\binom{n}{k}$ is the binomial coefficient. $y_{p}^{\text{emb}}$ is the token-weighted sentence embedding of topic $p$, and $y_{s}^{\text{emb}}$ is the token-weighted sentence embedding of all topics in subset $S$. For each combination of topic embeddings, I calculate the $R_{\text{oos}}^{2}$ accordingly. [^4] Shapley value requires an input for $X$ when it is missing. I assign a value of zero, treating the embedding as having all dimensions set to zero for model estimation. This approach fully distributes the total $R^{2}_{\mathrm{OOS}}$ among topics, revealing their relative importance in explaining market reactions.
@@ 3 Data and Empirical Results
This section presents the empirical analysis of analyst report information. I begin by examining the information content and decomposing it across different topics. I then quantify the strategic dollar value of analyst information. The results emphasize analysts’ income statement analyses following earning announcements.
@@ 3.1 Data
Sell-side analyst reports are obtained from the Mergent Investext Database. I download 223,091 analyst reports of S&P 100 firms from 2000 to 2023. The initial reports are in PDF format, and I use Adobe Acrobat to extract plain text from each report. The text is chunked into sentences. To get clean content of firm-relevant discussions, I fine-tune a BERT model to remove the boilerplate segments. [^5] Boilerplate segments disclose information about the brokerage firm and analyst research team and do not reflect analyst opinions. Following Li et al. (2024), I randomly sampled one report from each of the top 20 brokerage firms producing the largest number of analyst reports and manually labeled each sentence in those reports. The BERT model fine-tuned for boilerplate classification reached an accuracy score of 97.31% on a testing sample.
@@
The Mergent Investext Database provides information on analyst names and the release date of each report. To connect Investext with the I/B/E/S database, I first match the lead analyst name in reports to Analyst ID (AMASKCD) following the procedure of Li et al. (2024). I then match specific reports to the corresponding EPS forecasts, price targets, and recommendation files in the I/B/E/S database. Since multiple reports could be released between a forecast announcement and its revision, I follow Huang et al. (2014) and define the matching window for each report as 2 days before the announcement date to 2 days after the revision date. [^6] In I/B/E/S, each analyst forecast for a specific firm has two time stamps. The announcement date (ANNDATS) is the date the analyst announces the new forecast and revises the old one. The revision date (REVDATS) is the date the forecast becomes invalid or is revised. Reports are linked to I/B/E/S forecasts by matching analyst names, firm tickers, and the matching window. I also compile firm characteristics from the CRSP and Compustat databases. [^7] See Table A2 and Table A3 for detailed variable definitions and summary statistics. After the matching procedure and excluding analyst reports released on earnings announcement days, my final sample comprises 122,252 reports of S&P 100 stocks from 1,305 analysts across 140 unique brokerage firms.
@@
Table 1 presents the summary statistics of the analyst report data, including the number of reports, the number of unique brokerage firms, and the number of unique analysts across years and the Fama-French 12 industries. It also summarizes the average number of pages and tokens per report, which are important indicators of report lengths. Overall, the time trend in the number of analyst reports aligns with that reported in Bonini et al. (2023), and the average number of pages per report is consistent with that in Huang et al. (2014). [^8] The data reveals a peak in the number of analyst reports in 2013, followed by a subtle decline. This trend can be attributed to the regulatory impacts of the Dodd-Frank Wall Street Reform in the United States and the introduction of MiFID II in Europe (Bradshaw et al., 2017).
@@
The stock-level price impact is estimated using intraday data from the NYSE TAQ database, covering an out-of-sample period from January 2, 2015, to December 28, 2023. To ensure comparability over time, all prices are adjusted for inflation using the Consumer Price Index (CPI), with December 2020 as the base period.
@@
Given my goal to estimate the information value of analyst outputs, I need to control for other information available to investors. In Section 3.2.2, I compare textual information from earnings conference calls and subsequent analyst reports. Earnings conference call transcripts are sourced from the Seeking Alpha website, which has been collecting the transcripts since 2005999Available at https://seekingalpha.com/earnings/earnings-call-transcripts.. These transcripts provide comprehensive record of the calls, including executive presentations and manager-analyst interactions like Q&A sessions. I sample earnings conference call transcripts of S&P 100 constituent firms from 2005 to 2023, merging them with ticker and Report Date for Quarterly Earnings (RDQ) data from Compustat. The matching window between Seeking Alpha’s publishing date and RDQ is set to two weeks. This process yielded a final sample of 7,909 firm-quarter-level conference calls.
The set of numerical variables used as input in machine learning models includes:
@@
Firm-Level Measures: $Size$ is measured by market equity. The Book-to-Market Ratio ($BtoM$) compares the firm’s book value to its market value. Following Huang et al. (2018), the Earnings Suprise ($SUE$) is calculated as the actual earnings per share (EPS) minus the last consensus EPS forecast prior to the earnings announcement date (EAD), with $AbsSUE$ representing its absolute value. The $Miss$ variable is a binary indicator set to one if $SUE$ is positive. $TradingVolume$ is the stock’s trading volume on the latest earnings announcement days standardized by the number of shares outstanding. The Distance to Default ($DD$) variable is sourced from the NUS Credit Research Initiative (CRI). $Fluidity$ measures firms’ product market fluidity.
@@ 3.2 Information Content of Analyst Reports
In this section, I examine the information content of analysts’ quantitative and qualitative outputs by projecting contemporaneous cumulative abnormal returns on analyst information. Furthermore, I attribute the information content of analyst report text across topics using the Shapley value decomposition introduced in Section 2.4.
@@ 3.2.1 Information Content of Quantitative and Qualitative Information
Table 2 compares the information content embedded in numerical and textual measures, quantified with $R^{2}_{\mathrm{OOS}}$. The estimation model applied in this table is the Ridge regression specified in Equations 1 and 2. The objective variable is ${CAR}_{[-1,+1]}$ centered at the analyst report release date. For reports released on non-trading days, I adjust t=0 to the next trading day recorded in the CRSP database. I estimate separate Ridge regression models for distinct input configurations and employ an expanding window approach, constructing training sets from 2000 to the year preceding each target year.
@@
Panel A of the table delineates the performance of four types of input configurations: (1) forecast revisions, encompassing analyst revisions on recommendations, earnings forecasts, and target prices; (2) numerical measures, incorporating report-level, firm-level, industry-level, and macroeconomic indicators as detailed in Section 3.1; (3) textual embeddings, 5,120-dimensional full-context representations of each report; and (4) a hybrid input combining revision measures with textual embeddings.
@@
Column (1) reveals an $R^{2}_{\mathrm{OOS}}$ of 9.01% with forecast revision input, corroborating extant literature (Brav and Lehavy, 2003; Asquith et al., 2005; Bradshaw et al., 2013) on the price impact of analyst forecast revisions. The inclusion of additional numerical variables in column (2) yields only a marginal improvement, with the $R^{2}_{\mathrm{OOS}}$ reaching 9.08% using 17 numerical features. I compare the estimation with a zero benchmark and report the Diebold-Mariano test t-statistics in columns (2) and (4). The models with revision and numerical inputs significantly outperform a zero-estimation benchmark.
@@
Columns (5) and (6) elucidate the performance of Ridge models with text embeddings as input, showcasing a substantial $R^{2}_{\mathrm{OOS}}$ of 10.19% and a t-statistic of 8.20, underscoring the significant information content of the report text in explaining contemporaneous stock returns. The $R^{2}_{\mathrm{OOS}}$ exhibits relative stability across 2015-2023, with a notable exception during the 2020 pandemic recession. The DM test comparing revision-only and text-only inputs yields a t-statistic of 1.66, suggesting substantial information content in analyst report text compared with forecast revisions.
Sarkar and Vafa (2024) raise the look-ahead bias concern associated with pre-trained LLMs. Language model outputs may incorporate future information that should not be available. Since the LLMs I applied were trained on language data ending before 2023, the $R^{2}_{\mathrm{OOS}}$ in 2023 reflects true out-of-sample performance, confirming the robustness of the main findings during this period.
@@
In columns (7) and (8), I further investigate the model performance when combining both forecast revisions and text embeddings in the Ridge regression. The combination achieves an $R^{2}_{\mathrm{OOS}}$ of 12.28%, which represents a 3.27% improvement over the $R^{2}_{\mathrm{OOS}}$ of the revision-only model and a 2.09% improvement over the $R^{2}_{\mathrm{OOS}}$ of the text-only model. I report the DM t-statistic of the comparison between combined input with revision input and text input in columns (7)-(1) and (7)-(5). The overall t-statistics of 3.95 and 3.77 indicate a significant superiority in the combination of text data and revision measures in explaining contemporaneous $CAR$. The DM tests confirm that report text and forecast revision contain distinct information sets that are valued by investors.
@@
The explanatory power might be derived from quantitative information in the numbers mentioned in report text. I conduct robustness checks using content with numbers removed. To address the concern, I exclude numerical data from the written reports before generating text embeddings and report model performance in Table A4. The $R^{2}_{\mathrm{OOS}}$ of the number-free text embedding input model is 10.95%, significantly higher than that of quantitative revision measures (t-statistic = 2.40). The robustness check suggests that the superior performance of text-based models is not driven by the numbers embedded in text.
@@
LLMs exhibit improved language skills as the parameter size increases. I provide a quantitative comparison of LLMs with varying numbers of parameters, including the BERT model, OpenAI text-embedding-3-small model, and LLaMA-3 model. The LLaMA models have a substantially larger number of parameters compared to the BERT and OpenAI models. Panel B reports the $R^{2}_{\mathrm{OOS}}$ and DM t-stats of pairwise comparisons between alternative LLMs and the LLaMA-2 model. The input in Panel B is report text embeddings, comparable to column (3) in Panel A. A positive t-statistic indicates that the model outperforms the LLaMA-2 benchmark. Overall, the BERT and OpenAI models underperform the LLaMA-2 model. The LLaMA-3 model gives an $R^{2}_{\mathrm{OOS}}$ of 9.66%, which is slightly better but statistically insignificant compared to the LLaMA-2 model. The results in Table 2 suggest that the out-of-sample information content of analyst report text is fairly robust across different LLMs. Pairwise comparisons between LLMs also shed light on how the scale of the model influences performance capabilities.
@@
To test the robustness of the information content in these contemporaneous measures, I also use $CAR$ measures in 2-day and 5-day windows, as applied in Huang et al. (2014) and Asquith et al. (2005), respectively. Additionally, I examine the abnormal stock return on the days that reports are released. In Table A5, I find a similar magnitude in $R^{2}_{\mathrm{OOS}}$ for explaining $AR_{[0]}$, $CAR_{[0,+1]}$, and $CAR_{[-2,+2]}$, ranging from 6.92% to 7.57% for the out-of-sample period from 2015 to 2023.
@@
In Table A6, I present the information content of analyst report text across the Fama-French 12 industries. The $R^{2}_{\mathrm{OOS}}$ exceeds 10% in five industries: Shops, Other, Manufacturing, Chemicals, and Durables. Notably, the $R^{2}_{\mathrm{OOS}}$ and DM t-statistics are significantly lower in the Non-Durables, Telecommunications, Energy, and Utilities sectors. The industries with higher $R^{2}_{\mathrm{OOS}}$ typically involve complex products and services that benefit from detailed analyst insights, whereas those with lower $R^{2}_{\mathrm{OOS}}$ may operate in more stable or regulated environments where analyst reports add less incremental value. In Figure A3, I plot analyst information value by Fama-French 12 industries, revealing a pattern consistent with the information content results.
To assess the economic magnitude, I perform OLS regressions specified as follows: